# Loss of Control Prevention of an Agile Aircraft: Dynamic Command Saturation Approach

### v1

@@ 3.2 Control Allocation Design
The need for a control allocation algorithm due to over-actuation has already been discussed. Thus, to achieve the integrity of the flight control architecture, an optimization-based control allocation is utilized. The objective function represents the minimization of control effort, subject to the position limits. Specifically, the flight control allocation problem can be formulated as given in Eq. (7).
@@
where $\Phi(x_{0},u_{0})\in\mathbb{R}^{3\times m}$ is the Jacobian matrix at that instant, where $m$ is the number of control surfaces, and $\tau_{c}$ represents the commanded control moment coefficients. Additionally, the upper and lower bounds, represented by $u_{min}$ and $u_{max}$, significantly influence the derived control surface deflections. Furthermore, the Jacobian matrix is calculated at each time step using the central-difference method to capture the aerodynamic nonlinearity.
@@
As illustrated in Fig. 1, the pilot inputs $\alpha_{cmd}$, $\beta_{cmd}$, and $p_{cmd}$ are subjected to the incremental nonlinear dynamic inversion to derive the control moment coefficients. Afterwards, the control moment coefficient are transferred to the control allocation module to derive the corresponding control surface deflections while minimizing the control effort. All in all, the control surface deflections are injected into the nonlinear flight dynamics module for deriving new states, $x$. Notice that, since the inner-loop is also derived using incremental nonlinear dynamic inversion, the finalized control surface deflections are also in the incremental form. This implies that the control allocation is only responsible for distributing the necessary control moment coefficients on the control surfaces in an optimal manner.
4 Incremental Attainable Moment Set
The attainable moment set (AMS) of an aircraft is a geometrical representation (topologically defines a convex-hull) of the aerodynamic moments that can be generated by the aircraft’s control effectors at various flight conditions. Furthermore, an attainable moment set of an aircraft helps understanding the instantaneous control authority of the aircraft, i.e. controllability of the aircraft [31, 32]. Due to this benefit, the AMS will be used in this study to detect the possibility of loss of control. However, constructing the attainable moment set regarding only the actuator position limits implies that one assesses the attainable moment set of an aircraft, under any arbitrary circumstance, can observe the global achievable control moments. Instead, an alternative path is proposed to give a better insight about the instantaneous controllability during an agile maneuver, i.e. incremental attainable moment set.
The concept and utilization of the incremental attainable moment set (IAMS) is the same as the attainable moment set with a distinction: the boundary formulation. Not only the actuator deflection limits but also the actuator rate limits should be considered to understand whether the aircraft is capable of providing necessary control moment at that time step or not. Therefore, with a slight notion shift, the attainable moment set is reformulated to construct the incremental attainable moment set $\Delta\mathbb{A_{M}}$ as given in Eq. (8).
@@ 5 Loss of Control Prevention
The loss of control notion has been discussed in the context of controllability, and this study specifically addresses the prevention of departure. Furthermore, the prevention architecture is divided into two stages: detection and prevention, both of which are proposed in an online manner.
  • Detection: The IAMS serves as the detector for the initiation of the loss of control, indicating the available control authority. The flight control architecture derives the desired control moment vector, which is then compared to the instantaneous IAMS to determine whether the demand falls within the bounds of the available control authority. However, as a safety margin, the acquired instantaneous IAMS is shrunk by a factor of 30%, meaning that only 70% of the available control authority is allowed to be used. This is not a strict requirement but can be modified based on expectations (the greater the shrinkage factor, the more restrictive it is). The specified level is regarded as sufficiently satisfactory in this study, considering the trade-off between maneuverability and loss of control.
  • Prevention: A dynamic (online) command saturation approach is developed in order to prevent the loss of control (i.e. departure) due to the excessive or abrupt pilot inputs. In case of the detection of the violation of the control authority, the dynamic command saturation module is activated and the states of $\alpha$, $p$, $q$ and $r$ are restricted. The dynamic command saturation is designed based on the Lyapunov’s direct method, which will be discussed in the proceeding section.
Consequently, the primary objective is to prevent departure during extreme maneuvers by saturating the pilot commands in the event of a loss of control detection. In this way, executing the most agile maneuvers without entering into an upset condition is desired.
5.1 Lyapunov-based Dynamic Command Saturation Design
@@
Consequently, the stabilizing control moments is derived using Lyapunov’s direct method, and these regulator moments will be utilized to derive the pilot command saturation expressions. By using Euler’s equations of motion, the angular rate saturation can be established through the following relation given in Eq. (16).
@@
where the expression of $[q\hskip 4.26773ptr]_{sat}^{T}$ is derived in Eq.(18). However, only the angle of attack is saturated since the sideslip angle is desired to be zero for all flight circumstances. Consequently, through inductive reasoning, the expression for dynamic command saturation can be derived. By designing the stabilizing control moments, a correlation between them and the dynamic command saturation expressions can be established. Therefore, an online solution for the command saturation is proposed. The only necessity for implementing the proposed online architecture is accurate measurement and a high-fidelity, broad aerodynamic database. The remaining steps involve simple algebraic calculations, so the computational cost depends only on deriving the necessary aerodynamic coefficients. After all, for the sake of clarity, a flowchart of the proposed logic is presented in Fig. 2.
@@ 6.1 Abrupt & Agile Maneuver Evaluation
The preliminary assessments of the proposed architecture are presented with two distinct agile maneuvers [^5] For the simulation videos, please visit: https://youtu.be/I524wsm6EPk. The agility metrics of the commands are defined as follows: high amplitude control commands, abrupt application, and simultaneous application. Based on these metrics, angle of attack and roll rate inputs are applied abruptly, with high amplitude, and simultaneously. Furthermore, the static stability degradation has a significant impact on the loss of control; that is, the worst-case scenario is selected, and the center of gravity is located such that the aircraft is statically unstable (at %$\%35\bar{c}$). Subsequently, the effects of the proposed architecture are presented after demonstrating that the aircraft is not capable of performing the desired maneuver without the loss of control prevention architecture. The maneuver properties are detailed in Table 2.
@@ 6.1.1 Maneuver #1 Examination
Obviously, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space, i.e., the attainable moment set, at the instant of input reversals. The surpassing of controllability boundaries initiates a departure from stable flight, resulting in the occurrence of loss of control. Subsequent to stimulation of the loss of control, the control deflections grow irrationally and even reach saturation. Additionally, the discrepancy between the required and achieved control moment coefficients progressively worsens until the simulation is terminated due to numerical issues arises from the loss of control. However, the effects of the loss of control prevention architecture implementation are shown in Fig. 4.
@@
Furthermore, the dynamic variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 5.
@@
The attainable moment set history for both loss of control prevention is on and off is demonstrated in Fig. 6.
@@
Attainable moment set illustrations of the maneuver #1 with loss of control prevention indicates the arbitrary instantaneous controllability checks as stated previously, and it is observed that at $t_{3}$ the desired control moment vector violates the attainable moment set, i.e. red, while there is not loss of control prevention. Remember from Figure 3 that the violation stimulates the loss of control and the aircraft departures. However, in case of activation of the loss of control prevention architecture, the demanded control moment coefficient vector is still bounded by the attainable moment set; therefore, any loss of control is avoided. Consequently, the history of the control deflections with demanded moment satisfaction parameter and the 3D trajectory of the maneuver are depicted in Fig. 7 and Fig. 8 respectively.
@@
Obviously, without the loss of control prevention architecture, the aircraft is incapable of executing the desired maneuver. Eventually, the simulation demonstrates that the aircraft outperforms this maneuver compared to the simulation without the loss of control prevention. Additionally, it is worth highlighting that the maneuver is so agile that the load factor approaches the limit of 10g.
@@ 6.1.2 Maneuver #2 Examination
Similar to the previous instance, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space. Once again, surpassing the controllability boundaries initiates a departure from stable flight, resulting in the occurrence of loss of control. The effects of the loss of control prevention architecture implementation are depicted in Fig. 10.
@@
Furthermore, the dynamic variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 11.
@@
Ultimately, without the loss of control prevention architecture, the aircraft is unable to execute the desired maneuver. It is proven that the aircraft outperforms this maneuver compared to the simulation without the loss of control prevention. It is also noteworthy that the maneuver is so agile that the load factor approaches the limit of 15g.
@@ 6.2 Comparison Study: Scheduled State Limiter versus Lyapunov-based Command Saturation
Figure 15: The comparison for sea level altitude with various Mach number: the volume of the yellow shaded polygon is 0.9504uc whereas the volume of the red shaded polygon is 0.7111uc.
Figure 16: The comparison for 10000ft altitude with various Mach number: the volume of the yellow shaded polygon is 0.7907uc whereas the volume of the red shaded polygon is 0.5094uc.
The most eye-catching distinction of these two separate assessment’s properties is the reduction in the most agile and stable maneuver capability, which is highly related to the control authority alteration due to the available dynamic pressure. Furthermore, an observable distinction in achievable positive and negative roll rates exists because of the asymmetric properties of the aircraft in terms of inertia and engine.
However, perhaps the most significant interpretation of these assessments is that the Lyapunov-based command saturation approach allows for more agile maneuvers without compromising stable flight, which has been the main objective of the study. In terms of expansion, the Lyapunov-based command saturation approach increased the generated volume compared to the conventional approach by 33.66% for the sea level altitude assessment and by 55.21% for the 10000ft altitude assessment. This serves as concrete evidence that more agile and stable flight can be achieved using the proposed method. However, it is noteworthy that, even using the proposed method, not all combinations of angle of attack and roll rate maneuvers can be performed without losing control.
@@ 7 Conclusion & Future Works
In this study, an online loss of control prevention methodology is presented for over-actuated F-16 model through merging the Lyapunov’s direct method and inverse dynamics of the aircraft. The flight control architecture is established based on the design of flight control law using incremental nonlinear dynamic inversion and an optimization-based control allocation due to the over-actuation, whereas the pilot inputs are selected as the angle of attack ($\alpha$), sideslip angle ($\beta$), and roll rate ($p$). However, the challenge of the proposed methodology is not only to control the aircraft but also to ensure the flight safety by eliminating the destabilizing effects which has been discussed.
To address this critical issue, a structure with two components is developed: a controllability investigation module (using incremental attainable moment set) and a dynamic command saturation module (merging Lyapunov’s direct method and inverse dynamics of the aircraft). Through the proposed structure, the control commands are dynamically saturated during flight to prevent loss of control in the event of a detection of a loss of control possibility. Prior to applying dynamic saturation, a controllability assessment is conducted using an incremental attainable moment set, which considers actuator limitations such as position and rate saturation. This assessment determines whether the demanded control moment vector falls within the attainable set or not, whereas the allowable control authority is set at 70%, with a 30% safety margin. Violating this set indicates an excess of control authority, and in such cases, the Lyapunov-based dynamic command saturation is activated.
The effectiveness and performance of the proposed architecture are demonstrated through two distinct and challenging maneuvers. It is shown that the aircraft is unable to perform these maneuvers without the proposed loss of control prevention structure. However, by activating the loss of control prevention approach, stable maneuvers can be maintained. Furthermore, a comparison is made between the conventional state limiters and dynamic command saturation, revealing a significant expansion of the stable maneuverable volume using the proposed method (33.66% at the sea level and 55.21% at 10000ft). In conclusion, the proposed architecture extends the safe maneuvering envelope of the aircraft, enabling the execution of most abrupt and agile maneuvers without departing from stable flight.
Further studies are planned to examine the effects of the scheduled gain on the expansion of the stable maneuverable volume using an adaptive Lyapunov gain. Additionally, another aim is to explore the compatibility of the presented method with different baselines, such as commercial aviation or quadcopters. Hence, the plan includes working on the real-time implementation of various baselines and investigating scenarios involving a fault occurrence, utilizing either the presented methodology or an improved version, as a part of future work.

### v2

@@ 3.2 Control Allocation Design
The need for a control allocation algorithm due to over-actuation has already been discussed. Thus, to calculate the necessary control surface deflections, an optimization-based control allocation is utilized. The objective function represents the minimization of the demand on the control system, with control surfaces position limiting. Additionally, the demand on the control system is quantified using the square of the Euclidean 2-norm, denoted as $||u||^{2}_{2}$ or $u^{T}u$ in the vector form. Specifically, the flight control allocation problem can be formulated as given in Eq. (9).
@@
where $\Phi(x_{0},u_{0})\in\mathbb{R}^{3\times m}$ is the Jacobian matrix at that instant, where $m$ is the number of control surfaces, and $\tau_{c}$ represents the commanded control moment coefficients. Additionally, the upper and lower bounds of the actuator position limits, represented by $u_{min}$ and $u_{max}$, significantly influence the derived control surface deflections. Furthermore, the Jacobian matrix is computed at each time step using the central-difference method. This approach captures the aerodynamic nonlinearity through discretization, accommodating the nonlinear and complex aerodynamic characteristics.
@@
As illustrated in Fig. 1, the pilot commands $\alpha_{cmd}$, $\beta_{cmd}$, and $p_{cmd}$ are subjected to the incremental nonlinear dynamic inversion to derive the control moment coefficients. Afterwards, the control moment coefficient are transferred to the control allocation module to derive the corresponding control surface deflections while minimizing the control effort. All in all, the control surface deflections are injected into the nonlinear flight dynamics module for deriving new states, $x$. Notice that, since the inner-loop is also derived using incremental nonlinear dynamic inversion, the finalized control surface deflections are also in the incremental form. This implies that the control allocation is only responsible for distributing the necessary control moment coefficients on the control surfaces in an optimal manner, as the main objective is minimizing the control effort while satisfying the desired control moment coefficients previously described.
@@ 4 Loss of Control Prevention
The loss of control notion has been discussed in the context of controllability, and this study specifically addresses the prevention of departure. Furthermore, the architecture is divided into two stages: detection and prevention, both of which are proposed in an online manner.
  • Detection: The IAMS serves as the detector for the initiation of the loss of control, indicating the available control authority. The flight control architecture derives the desired control moment vector, which is then compared to the instantaneous IAMS to determine whether the demand falls within the bounds of the available control authority. However, as a safety margin, the acquired instantaneous IAMS is shrunk by a factor of 30%, meaning that only 70% of the available control authority is allowed to be used. This direct correlation between the IAMS and the available control authority is achievable, provided that the aerodynamic database is extensive and high-fidelity, accounting for nonlinearities and over-actuation effects as comprehensively as possible. Moreover, the utilized shrinkage factor is not a strict requirement but can be modified based on expectations (the greater the shrinkage factor, the more restrictive it is). The specified level is regarded as sufficiently satisfactory in this study, considering the trade-off between maneuverability and loss of control. Hence, the approach will be conservative when a moment cannot be instantaneously achieved, and the prevention mechanism in the controller saturates the command.
  • Prevention: An online command saturation approach is developed in order to prevent the loss of control (i.e. departure) due to the excessive or abrupt pilot inputs. In case of the detection of the violation of the control authority, the online command saturation module is activated and the states of $\alpha$, $p$, $q$ and $r$ are restricted. The online command saturation is designed based on the Lyapunov’s direct method, which will be discussed in the proceeding section.
Consequently, the primary objective is to prevent departure during extreme maneuvers by saturating the pilot commands in the event of a loss of control detection. In this way, executing maneuvers as agilely as the proposed strategy permits without entering into an upset condition is desired.
4.1 Lyapunov-based Online Command Saturation Design
@@
Consequently, the stabilizing control moments in Eq. (14) are derived using Lyapunov’s direct method, and these regulator moments will be utilized to derive the pilot command saturation expressions. By using Euler’s equations of motion, the angular rate saturation can be established through the following relation given in Eq. (16).
@@
where the expression of $[q\hskip 4.26773ptr]_{sat}^{T}$ is derived in Eq.(18). However, only the angle of attack command is intervened since the sideslip angle is commanded to be zero in all scenarios. This assumption is based on the expectation that a zero sideslip command can be tracked with only minor deviations. Furthermore, since Eq. (21) accounts for instantaneous states, the effects of a non-zero sideslip can be reflected to the yielded stabilizing angle of attack command value. Consequently, through inductive reasoning, the expression for online command saturation can be derived. By designing the stabilizing control moments, a correlation between them and the online command saturation expressions can be established. Therefore, an online solution for the command saturation is proposed. Implementing the proposed online architecture requires accurate real-time measurements and a high-fidelity aerodynamic database that covers a broad validity envelope, including post-stall conditions, with fine discretization and sufficient margins to account for the full range of the aircraft’s service envelope. The remaining steps involve simple algebraic calculations, so the computational cost depends only on deriving the necessary aerodynamic coefficients. After all, for the sake of clarity, a flowchart of the proposed logic is presented in Fig. 2.
@@ 5.1 Abrupt & Agile Maneuver Evaluation
The preliminary assessments of the proposed architecture are presented with two distinct agile maneuvers [^5] For the simulation videos, please visit: https://youtu.be/I524wsm6EPk. The agility of the aircraft is evaluated based on the following command metrics: (1) high amplitude control commands (commands are considered high amplitude if they involve changes greater than 30 for pitch rate and 90 for roll rate); (2) abrupt application (commands are considered abrupt if they change considerably in a manner similar to step-wise changes); (3) simultaneous application (commands are considered simultaneous if they are applied concurrently). Based on these metrics, angle of attack and roll rate inputs are applied abruptly, with high amplitude, and simultaneously. Furthermore, the static stability degradation has a significant impact on the loss of control; that is, the worst-case scenario is selected, and the center of gravity is located such that the aircraft is statically unstable ($x_{CG}$ = %$\%35\bar{c}$). Subsequently, the effects of the proposed architecture are presented after demonstrating that the aircraft is not capable of performing the desired maneuver without the loss of control prevention architecture. The flight conditions and the $x_{CG}$ position set for the simulations are detailed in Table 2.
@@ 5.1.1 Maneuver #1 Examination
Obviously, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space, i.e., the attainable moment set, at the instant of input reversals. The violation of controllability boundaries initiates a departure from stable flight, resulting in the occurrence of loss of control. Subsequent to stimulation of the loss of control, the control deflections grow irrationally and even reach saturation. Additionally, the discrepancy between the required and achieved control moment coefficients progressively worsens until the simulation is terminated due to numerical issues arises from the loss of control. However, the effects of the loss of control prevention architecture implementation are shown in Fig. 4.
@@
Furthermore, the variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 5.
@@
The history of the attainable moment set with loss of control prevention both enabled and disabled is shown in Fig. 6.
@@
Attainable moment set illustrations of the maneuver #1 with loss of control prevention indicates the arbitrary instantaneous controllability checks as stated previously, and it is observed that at $t_{3}$ the desired control moment vector violates the attainable moment set, i.e. red, while there is not loss of control prevention. Remember from Fig. 3 that the violation stimulates the loss of control and the aircraft departures. However, in case of activation of the loss of control prevention architecture, the demanded control moment coefficient vector is still bounded by the attainable moment set; therefore, any loss of control is avoided. Consequently, the history of the control deflections with demanded moment satisfaction parameter and the 3D trajectory of the maneuver are depicted in Fig. 7 and Fig. 8 respectively.
@@
The control surface deflection history in Fig.7 puts forth that the desired control moment coefficients are achieved with remarkable accuracy, while the magnitudes of the control surface deflections remain relatively modest, indicating that the control effort is minimized as much as possible. Obviously, without the loss of control prevention architecture, the aircraft is unable to execute the desired maneuver effectively. The simulation results demonstrate that with the loss of control prevention enabled, the aircraft successfully performs the maneuver, in contrast to the unsuccessful attempt without this feature. It is important to note that the aircraft exhibits such high agility during the maneuver that the load factor approaches the 10g limit.
@@ 5.1.2 Maneuver #2 Examination
Similar to the previous instance, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space. Once again, violation of the controllability boundaries initiates a departure from controlled flight. The effects of the loss of control prevention architecture implementation are depicted in Fig. 10.
@@
Furthermore, the variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 11.
@@
Ultimately, without the loss of control prevention architecture, the aircraft is unable to execute the desired maneuver. It is proven that the aircraft outperforms this maneuver compared to the simulation without the loss of control prevention. It is also noteworthy that the maneuver is so extreme that the load factor approaches the limit of 15g.
@@ 5.2 Comparison Study: Scheduled State Limiter versus Lyapunov-based Command Saturation
Figure 15: The comparison for sea level altitude: the yellow polygon’s volume is 0.9504uc, while the red polygon’s is 0.7111uc.
Figure 16: The comparison for 10000ft altitude: the yellow polygon’s volume is 0.7907uc, while the red polygon’s is 0.5094uc.
The most eye-catching distinction of these two separate assessment’s properties is the reduction in the most agile and stable maneuver capability, which is highly related to the control authority alteration due to the available dynamic pressure. Furthermore, an observable distinction in achievable positive and negative roll rates exists because of the asymmetric properties of the aircraft, such as inertia coupling and the gyroscopic effect due to the engine rotating masses.
However, perhaps the most significant interpretation of these assessments is that the Lyapunov-based command saturation approach allows for more abrupt/higher amplitude maneuvers without compromising stable flight, which has been the main objective of the study. In terms of expansion, the Lyapunov-based command saturation approach increased the generated volume compared to the conventional approach by 33.66% for the sea level altitude assessment and by 55.21% for the 10000ft altitude assessment. The choice of using the expansion in the stable maneuverable volume as the assessment metric is fundamentally rooted in the objective of enabling higher amplitude maneuvers in a stable manner with the proposed approach, which aligns with the primary aim of this study. Therefore, this assessment serves as concrete evidence that more abrupt/higher amplitude and stable flight can be achieved using the proposed method. However, it is noteworthy that, even using the proposed method, not all combinations of angle of attack and roll rate maneuvers can be performed without losing control, probably reasoning from the lack of considering the dynamic pressure and inertial couplings in the loss of control detection through IAMS; additionally, the choice of Lyapunov gain $K$, which needs a further attention to be tuned.
@@ 6 Conclusion & Future Works
This study presents an online loss of control prevention method for an over-actuated F-16 model by combining Lyapunov’s direct method with the aircraft’s inverse dynamics. The flight control architecture integrates incremental nonlinear dynamic inversion and optimization-based control allocation, with pilot inputs being angle of attack ($\alpha$), sideslip angle ($\beta$), and roll rate ($p$). The challenge lies not only in controlling the aircraft but also in ensuring flight safety by eliminating destabilizing effects. To address this critical issue, a structure with two components is developed: a controllability investigation module (using incremental attainable moment set) and a online command saturation module (merging Lyapunov’s direct method and inverse dynamics of the aircraft). Through the proposed structure, the control commands are saturated during flight to prevent loss of control in the event of a detection of a loss of control possibility. Prior to applying online saturation, a controllability assessment is conducted using an incremental attainable moment set, which considers actuator limitations such as position and rate saturation. This assessment determines whether the demanded control moment vector falls within the attainable set or not, whereas the allowable control authority is set at 70%, with a 30% safety margin. In accordance with the assumptions and the scope of the study, violating this set indicates an excess of control authority, and in such cases, the Lyapunov-based dynamic command saturation is activated. The proposed architecture’s effectiveness is demonstrated through two challenging maneuvers, showing that stable maneuvers can be maintained only with the loss of control prevention structure. A comparison with conventional state limiters reveals a significant expansion in the stable maneuverable volume (33.66% at sea level and 55.21% at 10000 ft). The architecture extends the aircraft’s safe maneuvering envelope, enabling abrupt and agile maneuvers without compromising stability. Further studies are planned to examine the effects of the scheduled gain on the expansion of the stable maneuverable volume using an adaptive Lyapunov gain. Additionally, another aim is to explore the compatibility of the presented method with different aircraft types, such as commercial aviation airplanes or quadcopters. Hence, the plan includes working on the real-time implementation of various baselines and investigating scenarios involving a fault occurrence, utilizing either the presented methodology or an improved version, as a part of future work.